RC Meeting and Runplan, Wednesday, August 29

From HallCWiki
Jump to: navigation, search
  • Status
    • Reasonable spot size achieved at 3H07A and B harps yesterday evening, early OWL shift.
    • From that point - worked on sending CW to Hall C. Dealing with BLMs issues, etc. in machine.
    • Dave M. takes over as RC.
  • Runplan
    • Wednesday SWING, Tuesday OWL and DAY
      • Beam optics tuning
      • Big BPM calibration/checkout test plan
      • After initial setup - SHMS steering plan will be executed (please call RC if not already on site). When MCC asks, ramp up magnets (one at a time) to settings for -8.5 GeV.
      • Beam checkout
      • For the following, see detailed run plan here: [1]
        • Detector and trigger checkout. Call experts before start (after hole-target is centered): Malace, Pooser, Berdnikov, Ambrose for detector checkout, Sawatzky, Malace, Pooser, Huber, Trotta for trigger checkout)
        • Calibration runs
      • Note: Dave G. is seeing some odd results when comparing the BPMs and harps using scans from OWL shift Aug. 29. It would be useful to do a quick BPM/harp calibration in the near future if there's time.